# Dual-branch Spatial-Temporal Self-supervised Representation for Enhanced Road Network Learning

### v1

@@ Introduction
The road network is a vital infrastructure in a city, reflecting connectivity and accessibility between road segments, which guide human mobility. With the development of smart traffic systems (azgomi2018brief; camero2019smart), the road network has become a core component supporting various smart city applications, such as traffic inference (yang2022robust; zou2023novel) and destination prediction (xue2013destination; zong2019trip). Road network representation learning (RNRL) aims to learn universal and low-dimensional vector representations for road networks to enhance performance on downstream tasks (NEURIPS2024_15cc8e4a). Consequently, RNRL has gained significant attention from researchers and is emerging as a powerful tool for spatial-temporal management.
Early efforts leverage the natural graph topology of road network to learn task-specific representations based on Graph Neural Networks (GNNs) (jepsen2019graph; li2020spatial; gharaee2021graph). While these methods demonstrate robust performance on their original tasks, their efficacy often diminishes when transferred to different downstream tasks. Recent research on RNRL methods is inspired by powerful self-supervised learning paradigms to explore various geospatial relationships between roads through reconstruction and contrastive learning tasks (wu2020learning; chen2021robust). Additionally, abundant trajectory data are utilized to characterize the dynamic relationships between roads, augmenting road representations with auxiliary information (mao2022jointly; schestakov2023road). Nonetheless, the neighborhood smoothing mechanism inherent in GNNs leads to suboptimal road representations due to the spatial-temporal dynamics of roads, which adversely affects applications in dynamic downstream tasks.
@@
Figure 1 illustrates the representation challenges of spatial heterogeneity and temporal dynamics for RNRL. significant high-order interactions. Firstly, spatial heterogeneity reveals that road similarities can be estimated from several aspects beyond geospatial distance. For example, as illustrated in Figure 1(a), In the blue track, the next hop $S_{b}$ and the $S_{c}$ after multiple hops from road $S_{a}$ have configurations similar to $S_{a}$. Expanding the view to the road network, $S_{d}$, which is farther from $S_{a}$, serves the same function as $S_{a}$ (i.e., residential). Hence, methods restricted by nearby roads may yield an incomplete understanding of distant roads. Secondly, temporal traffic dynamics at different hours serve as a crucial complement to road representations, which cannot be adequately captured by road networks alone. As shown in Figure 1(b), distinct road types exhibit markedly divergent traffic volumes temporally. Critically, dynamic patterns demonstrate pronounced weekday-weekend divergence even for identical road categories. Therefore, road representation necessitates the consideration of both spatial heterogeneity and temporal dynamics to effectively adapt to dynamic downstream tasks.
@@
Here, $m$ denotes the number of roads within the single trajectory. We assign weights based on hop distance. In this context, the smaller hop distance receives larger initial weights, and the longer hop distance receives smaller initial weights. This strategy emphasizes adjacent links while incorporating reachable distant links.
@@
We expect the generalized rather than specific representation for downstream tasks through self-supervised learning. Therefore, we treat the representations $Z_{\mathcal{G}}$ and $Z_{\mathcal{H}}$ as two views of the road network for contrastive learning. We optimize the following MI maximization objective across these views (Zhu:2021tu).
@@
The coarse-grained road utilization feature fails to integrate into the temporal traffic flow characteristics. To this end, we employ Transformer-based encoders to model fine-grained traffic dynamics. The joint optimization of the dynamic prediction and classification tasks enables comprehensive learning of 24-hour traffic evolution trends and patterns.
@@
We expected to capture the patterns and trends of temporal traffic dynamic sequences through two tasks. Intuitively, the dynamic prediction task leverages the history sequence to predict the value of the sequence at the next time step, thereby capturing the long-term trend of the dynamic sequence. The dynamic classification task requires the model to categorize the input sequence into two types: weekday traffic and weekend traffic to distinguish between different dynamic sequence patterns. The corresponding regression value and classification probability are obtained after the representation passes through the project head, and these form the loss for the dynamic sequence task:
@@
where $_{\mathcal{G}}, ZZ_{\mathcal{H}}$, and $Z_{\mathcal{D}}$ are the updated spatial and temporal representations derived from the graph, hypergraph, and sequence encoder. $Z\in\mathbb{R}^{N\times d}$ is the final representation matrix of roads. The fusion methods we explored comprise concatenation along the last dimension, direct summation, and integration of representations via a gating mechanism. In this work, the first is employed for representation fusion.
@@
We used datasets from three real-world cities: Beijing, Porto, and Xi’an. The initial trajectory data consisted of GPS points collected by vehicles. We employed a map-matching algorithm (yang2018fast) to map it to the road sequence trajectory defined in the previous section. Please refer to Appendix C.1 for more details.
@@
Ours $\boldsymbol{0.7288}$ $\boldsymbol{0.8213}$ $\boldsymbol{236.6965}$ $\boldsymbol{363.2039}$ $\boldsymbol{0.6766}$ $\boldsymbol{0.8101}$ $\boldsymbol{101.4223}$ $\boldsymbol{143.6598}$ $\boldsymbol{0.8335}$ $\boldsymbol{0.8950}$ $\boldsymbol{202.8479}$ $\boldsymbol{307.7553}$
@@
Ours $\boldsymbol{2.4595}$ $\boldsymbol{3.2557}$ $\boldsymbol{3.4259}$ $\boldsymbol{4.5538}$ $\boldsymbol{4.4987}$ $\boldsymbol{5.6557}$
@@
Table 3: The results of zero-shot learning.
@@
Road networks are inherently structured as graphs. Most existing approaches draw inspiration from graph representation learning (zheng2022transition; zheng2023temporal) and employ graph neural networks (GNNs) for modeling. Numerous GNN-based methods  (wu2020learning; pei2020geom; zhang2023road; cao2025holistic) develop specialized graph encoders to capture the spatial topology of road networks. As trajectory data proliferates, trajectory-enhanced methods  (mao2022jointly; schestakov2023road; jiang2023self) extract movement and transfer patterns to advance road network representation learning. However, these methods inadequately model higher-order information within functional and movement. Our method integrates spatial semantic information from the perspectives of transfer mode and functional structure while incorporating travel traffic dynamics.
@@
Early graph learning (perozzi2014deepwalk; grover2016node2vec) focuses on the topology of graphs and learns the shallow representations of graphs. With the development of deep learning, graph convolutional network (kipf2016semi) and graph attention network (velivckovic2017graph) achieve significant success in graph learning by aggregating features in different ways. Temporal graph convolutional network considers dynamic changes in graphs (zhao2019t) later. Limited by the adjacency in graphs, these models struggle to represent higher-order relationships between nodes. Therefore, hypergraph learning is employed to model higher-order relationships among two or more entities. Hypergraph Neural Networks (feng2019hypergraph; gao2022hgnn+) adopt different message-passing mechanisms to learn hypergraph representations. However, neither graphs nor hypergraphs can be directly applied to road networks that require tailored designs due to their complex spatiotemporal characteristics.
@@ Acknowledgments
This work is supported in part by Zhejiang Provincial Natural Science Foundation of China (Grant No. LMS25F020012), the Hangzhou Joint Fund of the Zhejiang Provincial Natural Science Foundation of China under Grant No.LHZSD24F020001, Zhejiang Province High-Level Talents Special Support Program ”Leading Talent of Technological Innovation of Ten-Thousands Talents Program” (No.2022R52046), the Fundamental Research Funds for the Central Universities (2021FZZX001-23), and the advanced computing resources provided by the Supercomputing Center of Hangzhou City University.

### v2

@@ Introduction
The road network is a vital infrastructure in a city, reflecting connectivity and accessibility between road segments, which guide human mobility. With the development of smart traffic systems (azgomi2018brief; camero2019smart; wang2024cola; wang2024STAR), the road network has become a core component supporting various smart city applications, such as traffic inference (yang2022robust; zou2023novel) and destination prediction (xue2013destination; zong2019trip). Road network representation learning (RNRL) aims to learn universal and low-dimensional vector representations for road networks to enhance performance on downstream tasks (NEURIPS2024_15cc8e4a). Consequently, RNRL has gained significant attention from researchers and is emerging as a powerful tool for spatial-temporal management.
Early efforts leverage the natural graph topology of the road network to learn task-specific representations based on Graph Neural Networks (GNNs) (jepsen2019graph; li2020spatial; gharaee2021graph). While these methods demonstrate robust performance on their original tasks, their efficacy often diminishes when transferred to different downstream tasks. Recent research on RNRL methods is inspired by powerful self-supervised learning paradigms to explore various geospatial relationships between roads through reconstruction and contrastive learning tasks (wu2020learning; chen2021robust). Additionally, abundant trajectory data are utilized to characterize the dynamic relationships between roads, augmenting road representations with auxiliary information (mao2022jointly; schestakov2023road). Nonetheless, the neighborhood smoothing mechanism inherent in GNNs leads to suboptimal road representations due to the spatial-temporal dynamics of roads, which adversely affects applications in dynamic downstream tasks.
@@
Figure 1 illustrates the representation challenges of spatial heterogeneity and temporal dynamics for RNRL. Firstly, spatial heterogeneity reveals that road similarities can be estimated from several aspects beyond geospatial distance. For example, as illustrated in Figure 1(a), in the blue track, the next hop $S_{b}$ and $S_{c}$ after multiple hops from road $S_{a}$ have configurations similar to $S_{a}$. Extending the analysis to the broader road network shows that $S_{d}$, although spatially farther from $S_{a}$, serves a similar function (i.e., residential). Hence, methods restricted by nearby roads may yield an incomplete understanding of distant roads. Secondly, temporal traffic dynamics at different hours serve as a crucial complement to road representations, which cannot be adequately captured by road networks alone. As shown in Figure 1(b), distinct road types exhibit markedly divergent traffic volumes temporally. Critically, dynamic patterns demonstrate pronounced weekday-weekend divergence even for identical road categories. Therefore, road representation necessitates the consideration of both spatial heterogeneity and temporal dynamics to effectively adapt to dynamic downstream tasks.
@@
Here, $m$ denotes the number of roads within the single trajectory. The weights are assigned based on hop distance. In this context, the smaller hop distance receives larger initial weights, and the longer hop distance receives smaller initial weights. This strategy emphasizes adjacent links while incorporating reachable distant links.
@@
We pursue the generalized rather than specific representation for downstream tasks through self-supervised learning. Therefore, we treat the representations $Z_{\mathcal{G}}$ and $Z_{\mathcal{H}}$ as two views of the road network for contrastive learning. We optimize the following MI maximization objective across these views (Zhu:2021tu).
@@
The coarse-grained road utilization feature fails to integrate into the temporal traffic flow characteristics. To this end, we employ a Transformer-based encoder to model fine-grained traffic dynamics. The joint optimization of the dynamic prediction and classification tasks enables comprehensive learning of 24-hour traffic evolution trends and patterns.
@@
We aim to capture the patterns and trends of temporal traffic dynamic sequences through two tasks. Intuitively, the dynamic prediction task leverages the history sequence to predict the value of the sequence at the next time step, thereby capturing the long-term trend of the dynamic sequence. The dynamic classification task requires the model to categorize the input sequence into two types: weekday traffic and weekend traffic to distinguish between different dynamic sequence patterns. The corresponding regression value and classification probability are obtained after the representation passes through the project head, and these form the loss for the dynamic sequence task:
@@
where $_{\mathcal{G}}, ZZ_{\mathcal{H}}$, and $Z_{\mathcal{D}}$ are the updated spatial and temporal representations derived from the graph, hypergraph, and sequence encoder. $Z\in\mathbb{R}^{N\times d}$ is the final representation matrix of roads. The fusion methods we explore comprise concatenation along the last dimension, direct summation, and integration of representations via a gating mechanism. In this work, the first is employed for representation fusion.
@@
We use datasets from three real-world cities: Beijing, Porto, and Xi’an. The initial trajectory data consisted of GPS points collected by vehicles. We employ a map-matching algorithm (yang2018fast) to map it to the road sequence trajectory defined in the previous section. Please refer to Appendix C.1 for more details.
@@
Ours 0.7288 0.8213 236.6965 363.2039 0.6766 0.8101 101.4223 143.6598 0.8335 0.8950 202.8479 307.7553
@@
Ours 2.4595 3.2557 3.4259 4.5538 4.4987 5.6557
@@
Table 3: Zero-shot learning results (Beijing → Porto).
@@
Road networks are inherently structured as graphs. Most existing approaches draw inspiration from graph representation learning (zheng2022transition; zheng2023temporal; wang2024stega; yanggnns) and employ graph neural networks (GNNs) for modeling. Numerous GNN-based methods  (wu2020learning; pei2020geom; zhang2023road; cao2025holistic) develop specialized graph encoders to capture the spatial topology of road networks. As trajectory data proliferates, trajectory-enhanced methods  (mao2022jointly; schestakov2023road) extract movement and transfer patterns to advance road network representation learning. However, these methods inadequately model higher-order information within functional and movement. Our method integrates spatial semantic information from the perspectives of transfer mode and functional structure while incorporating travel traffic dynamics.
@@
Early graph learning (perozzi2014deepwalk; grover2016node2vec) focuses on the topology of graphs and learns the shallow representations of graphs. With the development of deep learning, graph convolutional network (kipf2016semi) and graph attention network (velivckovic2017graph) achieve significant success in graph learning by aggregating features in different ways. Temporal graph convolutional network considers dynamic changes in graphs (zhao2019t; wang2024unveiling) later. Limited by the adjacency in graphs, these models struggle to represent higher-order relationships between nodes. Therefore, hypergraph learning is employed to model higher-order relationships among two or more entities. Hypergraph Neural Networks (feng2019hypergraph; gao2022hgnn+) adopt different message-passing mechanisms to learn hypergraph representations. However, neither graphs nor hypergraphs can be directly applied to road networks that require tailored designs due to their complex spatiotemporal characteristics.
@@ Acknowledgments
This work is supported by the Starry Night Science Fund of Zhejiang University Shanghai Institute for Advanced Study (Grant No. SN-ZJU-SIAS-001), Zhejiang Provincial Natural Science Foundation of China (Grant No. LMS25F020012), the Hangzhou Joint Fund of the Zhejiang Provincial Natural Science Foundation of China under Grant No.LHZSD24F020001, Zhejiang Province High-Level Talents Special Support Program “Leading Talent of Technological Innovation of Ten-Thousands Talents Program” (No.2022R52046), the Fundamental Research Funds for the Central Universities (No.226-2024-00058), and the advanced computing resources provided by the Supercomputing Center of Hangzhou City University.